# Ahmed Heakl♠♣111 Equal Contributions , Abdullah Sohail111 Equal Contributions , Mukul Ranjan111 Equal Contributions , Rania Hossam111 Equal Contributions , Ghazi Ahmed Mohamed El-Geish , Omar Maher , Zhiqiang Shen , Fahad Khan♠♡ , Salman Khan♠♢

### v1

@@ Abstract
With the growing adoption of Retrieval-Augmented Generation (RAG) in document processing, robust text recognition has become increasingly critical for knowledge extraction. While OCR (Optical Character Recognition) for English and other languages benefits from large datasets and well-established benchmarks, Arabic OCR faces unique challenges due to its cursive script, right-to-left text flow, and complex typographic and calligraphic features. We present KITAB-Bench, a comprehensive Arabic OCR benchmark that fills the gaps in current evaluation systems. Our benchmark comprises 8,809 samples across 9 major domains and 36 sub-domains, encompassing diverse document types including handwritten text, structured tables, and specialized coverage of 21 chart types for business intelligence. Our findings show that modern vision-language models (such as GPT-4, Gemini, and Qwen) outperform traditional OCR approaches (like EasyOCR, PaddleOCR, and Surya) by an average of $60\%$ in Character Error Rate (CER). Furthermore, we highlight significant limitations of current Arabic OCR models, particularly in PDF-to-Markdown conversion, where the best model Gemini-2.0-Flash achieves only $65\%$ accuracy. This underscores the challenges in accurately recognizing Arabic text, including issues with complex fonts, numeral recognition errors, word elongation, and table structure detection. This work establishes a rigorous evaluation framework that can drive improvements in Arabic document analysis methods and bridge the performance gap with English OCR technologies.
@@
Ahmed Heakl♠♣111 Equal Contributions , Abdullah Sohail111 Equal Contributions , Mukul Ranjan111 Equal Contributions , Rania Hossam111 Equal Contributions , Ghazi Ahmed Mohamed El-Geish , Omar Maher , Zhiqiang Shen , Fahad Khan♠♡ , Salman Khan♠♢ MBZUAI     Monta AI  Linköping University     Australian National University {ahmed.heakl,mabdullah.sohail,mukul.ranjan,salman.khan}@mbzuai.ac.ae {geish,omar}@monta.ai https://mbzuai-oryx.github.io/KITAB-Bench/
@@ 4.1 Evaluation Frameworks and Metrics
Chart-to-Dataframe: This task evaluates extracting structured data from Arabic charts into machine-readable dataframes. Systems must accurately parse numerical values, text labels, and preserve data relationships across chart types (bar, line, pie). We use the Structuring Chart-oriented Representation Metric (SCRM) Xia et al. (2024)—which combines type recognition, topic understanding, and structural numerical fidelity (see Appendix D)—and also propose our own CharTeX (Chart Extraction Score) metric. CharTeX combines the chrF scores for chart type and topic with the jaccord index for the dataframe, using fuzzy matching (80% threshold) when columns do not exactly align.
@@
Diagram-to-JSON: This task evaluates the conversion of Arabic flowcharts and technical diagrams into JSON while preserving semantic relationships and technical specifications. We propose CODM (Code-Oriented Diagram Metric), extending SCRM Xia et al. (2024), with the same fomulation as in Eq 3. More detail about this metric is provided in Appendix E.
@@ 4.2 Experimental Setup
We implement our evaluation pipeline with careful consideration of hyperparameters for different metrics. All experiments use NVIDIA A100 GPUs. For VLMs, we use their official implementations or API endpoints. Traditional OCR systems are evaluated using pre-trained models provided by the frameworks. For PDF-to-Markdown evaluation metric MARS 1, we choose $\alpha=0.5$ and $\alpha=0.5$ and $\beta=0.2$ for Diagram-to-JSON evaluation metric CODM. We average the results over multiple runs, with performance comparisons shown in different tables [3, 6, 5, 7, and 4].
@@ 5.3 Tables, OCR, and PDF-to-Markdown
Across table extraction tasks (Table 5), closed-source models maintain a clear advantage, with GPT-4o achieving 85.76% TEDS and 66.36% Jaccard scores. Among open-source models, AIN (75.94% TEDS) significantly outperforms Qwen variants, while specialized frameworks like Surya achieve competitive results (70.42% Jaccard) through targeted pipelines. In OCR evaluation (Table 6), Gemini-2.0-Flash leads with the lowest error rates (CER: 0.13, WER: 0.32). Notably, AIN matches this performance level (WER: 0.28), while traditional OCR frameworks like EasyOCR and Tesseract show moderate performance (CER: 0.58, 0.54). The significant performance drop in Paddle (CER: 0.79) and Surya (CER: 4.95) highlights the challenges in developing robust OCR systems.
@@ 7 Limitations and Future Directions
Despite its contributions, this benchmark has limitations. While it covers diverse Arabic document types, it lacks full representation of historical manuscripts and low-resource dialects. Future work should expand to include these, along with scanned records from government, academic, and financial institutions.
Another key limitation is in table and chart recognition, where OCR models struggle with structure preservation, header detection, and merged cell parsing. Though our benchmark introduces challenges in these areas, further refinements are needed for robust multimodal OCR capable of jointly processing text, tables, and figures. Future advancements should focus on dataset expansion, novel evaluation metrics, deep learning refinements, and cross-lingual OCR innovations to enhance Arabic VLMs.
@@ Appendix C Data Analysis
Our data generation pipeline (Figure 4) enabled the creation of 1,502 synthetic samples (576 charts, 422 diagrams, 456 tables). The pipeline’s human validation phase rejected 18% of initial outputs due to RTL formatting errors or semantic inconsistencies. As shown in Figure 5 and 6, domain-specific prompts ensured adherence to Arabic linguistic conventions during LLM-assisted generation. The final dataset exhibits balanced representation across:
  • Font Styles: 21 Arabic calligraphic styles
  • Document Types: 36 sub-domains including financial reports and technical manuals
  • Structural Complexity: 43% of tables contain merged cells; 29% of charts use dual-axis configurations
Appendix D Tasks Models and Metrics
@@
Appendix E SCRM and CODM
The Structuring Chart-oriented Representation Metric (SCRM) evaluates chart understanding through three components:
@@ Appendix E SCRM and CODM
where $J_{type}$, and $J_{topic}$ are chrF scores, and $J_{data}$ measures JSON structural similarity.
The Code-Oriented Diagram Metric (CODM) extends SCRM for flowcharts and technical diagrams:
$\text{CODM}=0.5J_{\text{topology}}+0.5J_{\text{semantics}}$ (5)
assessing both node-edge relationships and semantic labels. As shown in Figure 5 and 6, domain-specific prompts guided model responses for metric calculation. For instance, sequence diagrams required strict adherence to Arabic UML notation standards during evaluation.
@@
Dataset Size GPT-4o GPT-4o-mini Gemini-2.0-Flash Qwen2-VL
CER WER CER WER CER WER CER WER
PATS 500 0.23 0.30 0.53 0.71 0.01 0.02 1.02 1.02
SythenAR 500 0.09 0.20 0.14 0.32 0.07 0.17 0.59 1.13
HistoryAr 200 0.51 0.82 0.67 0.96 0.28 0.64 3.46 2.86
HistoricalBooks 10 0.41 0.76 0.59 0.88 0.05 0.22 1.90 2.16
Khatt 200 0.45 0.74 0.64 0.91 0.19 0.45 1.12 5.04
Adab 200 0.30 0.73 0.35 0.83 0.19 0.56 0.63 1.08
Muharaf 200 0.56 0.90 0.63 0.94 0.33 0.69 3.57 2.87
OnlineKhatt 200 0.29 0.63 0.41 0.76 0.17 0.44 1.30 2.01
ISI-PPT 500 0.08 0.18 0.15 0.31 0.06 0.15 1.03 1.06
ArabicOCR 50 0.06 0.26 0.16 0.46 0.00 0.02 1.25 1.50
Hindawi 200 0.34 0.56 0.48 0.71 0.01 0.04 1.82 2.05
EvArest 800 0.20 0.38 0.25 0.51 0.18 0.36 0.41 0.95
3,760 0.31 0.55 0.43 0.71 0.13 0.32 1.48 1.20
Dataset Size Qwen2.5-VL AIN Tesseract Surya
CER WER CER WER CER WER CER WER
PATS 500 0.26 0.36 0.00 0.00 0.14 0.28 4.66 4.67
SythenAR 500 0.21 0.40 0.04 0.16 0.31 0.72 4.82 7.90
HistoryAr 200 0.47 0.83 0.26 0.54 0.72 1.26 10.32 12.78
HistoricalBooks 10 0.33 0.72 0.84 0.88 0.74 0.99 6.81 6.30
Khatt 200 0.07 0.22 0.61 1.12 0.67 1.06 4.25 3.77
Adab 200 0.00 0.01 1.00 1.00 1.00 1.14 7.28 8.71
Muharaf 200 0.61 0.96 0.38 0.54 0.77 1.22 6.19 7.48
OnlineKhatt 200 0.36 0.70 0.03 0.12 0.59 1.20 6.71 6.95
ISI-PPT 500 0.36 0.54 0.52 0.53 0.31 0.64 4.25 3.77
ArabicOCR 50 1.00 1.00 0.01 0.01 0.01 0.01 2.75 3.58
Hindawi 200 1.00 1.00 0.11 0.15 0.31 0.72 0.15 0.20
EvArest 800 0.19 0.36 0.30 0.32 0.85 1.02 5.91 3.86
3,760 0.28 0.54 0.20 0.58 0.89 0.79 4.95 5.61

### v2

# Ahmed Heakl♠♣111 Equal Contributions , Abdullah Sohail111 Equal Contributions , Mukul Ranjan111 Equal Contributions , Rania Hossam111 Equal Contributions Ghazi Ahmad , Mohamed El-Geish , Omar Maher , Zhiqiang Shen Fahad Khan♠♡ , Salman Khan♠♢

@@ Abstract
With the growing adoption of Retrieval-Augmented Generation (RAG) in document processing, robust text recognition has become increasingly critical for knowledge extraction. While OCR (Optical Character Recognition) for English and other languages benefits from large datasets and well-established benchmarks, Arabic OCR faces unique challenges due to its cursive script, right-to-left text flow, and complex typographic and calligraphic features. We present KITAB-Bench, a comprehensive Arabic OCR benchmark that fills the gaps in current evaluation systems. Our benchmark comprises 8,809 samples across 9 major domains and 36 subdomains, encompassing diverse document types including handwritten text, structured tables, and specialized coverage of 21 chart types for business intelligence. Our findings show that modern vision language models (such as GPT-4o, Gemini, and Qwen) outperform traditional OCR approaches (such as EasyOCR, PaddleOCR, and Surya) by an average of $60\%$ in the character error rate (CER). Furthermore, we highlight significant limitations of current Arabic OCR models, particularly in PDF-to-Markdown conversion, where the best model Gemini-2.0-Flash achieves only $65\%$ accuracy. This underscores the challenges of accurately recognizing Arabic text, including issues with complex fonts, numeral recognition errors, word elongation, and table structure detection. This work establishes a rigorous evaluation framework that can drive improvements in Arabic document analysis methods and bridge the performance gap with English OCR technologies.
@@
Ahmed Heakl♠♣111 Equal Contributions , Abdullah Sohail111 Equal Contributions , Mukul Ranjan111 Equal Contributions , Rania Hossam111 Equal Contributions Ghazi Ahmad , Mohamed El-Geish , Omar Maher , Zhiqiang Shen Fahad Khan♠♡ , Salman Khan♠♢ MBZUAI   Monta AI   Linköping University  Australian National University {ahmed.heakl,mabdullah.sohail,mukul.ranjan,salman.khan}@mbzuai.ac.ae {geish,omar}@monta.ai https://mbzuai-oryx.github.io/KITAB-Bench/
@@ 4.1 Evaluation Frameworks and Metrics
Chart-to-Dataframe: This task evaluates extracting structured data from Arabic charts into machine-readable dataframes. Systems must accurately parse numerical values, text labels, and preserve data relationships across chart types (bar, line, pie). We use the Structuring Chart-oriented Representation Metric (SCRM) Xia et al. (2024)—which combines type recognition, topic understanding, and structural numerical fidelity (see Appendix D.1)—and also propose our own CharTeX (Chart Extraction Score) metric. CharTeX combines the chrF scores for chart type and topic with the jaccord index for the dataframe, using fuzzy matching (80% threshold) when columns do not exactly align.
@@
Diagram-to-JSON: This task evaluates the conversion of Arabic flowcharts and technical diagrams into JSON while preserving semantic relationships and technical specifications. We propose CODM (Code-Oriented Diagram Metric), extending SCRM Xia et al. (2024), with the same fomulation as in Eq 3. More detail about this metric is provided in Appendix D.1.
@@ 4.2 Experimental Setup
We implement our evaluation pipeline with careful consideration of hyperparameters for different metrics. All experiments use NVIDIA A100 GPUs. For VLMs, we use their official implementations or API endpoints. Traditional OCR systems are evaluated using pre-trained models provided by the frameworks. For PDF-to-Markdown evaluation metric MARS 1, we choose $\alpha=0.5$ and $\alpha=0.5$ and $\beta=0.2$ for Diagram-to-JSON evaluation metric CODM. We average the results over multiple runs, with performance comparisons shown in different tables (Table 3, 4, 5, 6, and 7).
@@ 5.3 Tables, OCR, and PDF-to-Markdown
Across table extraction tasks (Table 5), closed-source models maintain a clear advantage, with GPT-4o achieving 85.76% TEDS and 66.36% Jaccard scores. Among open-source models, AIN (75.94% TEDS) significantly outperforms Qwen variants, while specialized frameworks like Surya achieve competitive results (70.42% Jaccard) through targeted pipelines.
For OCR tasks, we evaluated GPT-4o  Hurst et al. (2024), Gemini-2.0-Flash  Google DeepMind (2025), Azure OCR Microsoft (2024) in closed model; Qaari NAMAA-Space (2025), Gemma3 Team et al. (2025), ArabicNagout Rashad (2024) and AIN Heakl et al. (2025) in open source models and Tesseract Smith (2007), EasyOCR JaidedAI (2020), PaddleOCR Li et al. (2022) and SuryaOCR Paruchuri (2024b) in frameworks (Table 6). Gemini-2.0-Flash leads with the lowest error rates (CER: 0.13, WER: 0.32). Notably, AIN matches this performance level (WER: 0.28), while traditional OCR frameworks like EasyOCR and Tesseract show moderate performance (CER: 0.58, 0.54). The significant performance drop in Paddle (CER: 0.79) and Surya (CER: 4.95) highlights the challenges in developing robust OCR systems.
@@
Figure 5: ChrF by model on Arabic text variations
@@
5.4 Performance on Challenging Cases
To evaluate model performance across different complexities of Arabic texts, we manually selected 104 samples representing four challenging categories: font variations, diacritics, text elongations, and tilted text. The ChrF score comparison (Figure 5) reveals distinct performance patterns across models, with GPT-4o demonstrating superior font handling (30.2) and leading in challenging tilted text recognition (13.1), while Azure OCR excels remarkably in diacritics recognition (40.5) and text elongations (35.2), indicating specialized Arabic script optimizations. The overall performance analysis shows GPT-4o leading at 26.0 average ChrF score, followed closely by Azure (26.3), Qwen2.5-VL-7B (25.3), and Gemini-2.0-Flash (24.7), while traditional OCR systems struggle significantly with Tesseract particularly challenged by diacritics (15.3) and tilted text (5.9). This analysis reveals that no single model excels across all Arabic text complexities, with specialized systems like Azure demonstrating domain-specific strengths in diacritics and elongation handling, while modern VLMs show more consistent performance but struggle with orientation variations, underscoring the need for Arabic-specific optimizations and highlighting the substantial performance gap between modern VLMs and traditional OCR approaches.
5.5 Model Performance across Chart Types
Figure 6: ChartEx results across different charts type.
The CharTeX evaluation across 16 different chart types reveals significant performance variations based on chart complexity and structural characteristics (Figure 6). Gemini-2.0-Flash demonstrates superior performance across most chart types, particularly excelling in simple geometric charts like Pie (72.4), Dot (72.7), and Bar Charts (69.2), while complex statistical visualizations like Violin Plots (32.3) and Box Plots (22.5) present significant challenges for all models. Simple chart types with clear boundaries consistently achieve higher scores across all models, with grouped and stacked bar charts showing intermediate performance levels around 40-50, indicating that while structural complexity affects extraction accuracy, the familiarity of bar chart formats provides some resilience. This pattern suggests that Arabic chart understanding faces particular difficulties with charts requiring statistical interpretation and continuous data representation, highlighting that current models perform best on charts with discrete, clearly separated data elements rather than continuous or overlapping visual representations.
@@ 7 Limitations and Future Directions
Despite its strengths, KITAB-Bench lacks coverage of low-resource dialects and institutional scans such as historical, governmental, and financial records. Future work should address OCR limitations in structural fidelity for tables and charts through richer datasets, refined metrics, and cross-lingual deep learning methods to enable robust and generalizable Arabic multimodal OCR. Moreover, current models often fail to generalize across domains and layouts, emphasizing the need for adaptable architectures and domain-specific fine-tuning.
@@ Appendix C Data Analysis
Our data generation pipeline (Figure 4) produced 1,502 high-quality synthetic samples - comprising 576 graphs, 422 diagrams, and 456 tables, through LLM-assisted generation guided by domain-specific instructions (Figures 7 and 8) that ensured alignment with Arabic linguistic norms. During the human validation phase, 18% of initial outputs were discarded due to issues like right-to-left formatting errors and semantic inconsistencies. The resulting dataset offers diverse and balanced coverage, featuring 21 Arabic calligraphic styles, 36 sub-domains spanning financial reports to technical manuals, and complex structures such as merged cells in 43% of tables and dual-axis configurations in 29% of charts.
Appendix D Evaluation Metrics
D.1 Tasks Models and Metrics
@@
D.2 Structuring Chart-oriented Representation Metric (SCRM)
The Structuring Chart-oriented Representation Metric (SCRM) evaluates chart understanding through three weighted components:
@@ D.2 Structuring Chart-oriented Representation Metric (SCRM)
where $J_{\text{type}}$ measures chart type recognition accuracy using Edit Distance, $J_{\text{topic}}$ evaluates chart topic identification using Edit Distance, and $J_{\text{data}}$ measures Mean Relative Error with and Error Thrsholding criteria.
For entity comparison in $J_{\text{type}}$ and $J_{\text{topic}}$, we use the chrF character-based metric which captures partial matches effectively. For data comparison, value similarity is computed using relative error:
$e(p,q)=\frac{|\text{Value}_{\text{pred}}^{p}-\text{Value}_{\text{GT}}^{q}|}{\text{Value}_{\text{GT}}^{q}}$
D.3 Chart Extraction Score (CharTeX)
To evaluate chart data extraction quality, we propose CharTeX (Chart Extraction Score), which combines character-level text similarity with structural data assessment:
$\text{CharTeX}=\alpha J_{\text{type}}+\beta J_{\text{topic}}+(1-\alpha-\beta)J_{\text{data}}$ (5)
Where $\alpha=0.05$ and $\beta=0.10$ in our implementation, reflecting the relative importance of each component where $J_{\text{type}}$ evaluates chart type recognition using chrF score (5% weight), $J_{\text{topic}}$ assesses topic identification using chrF score (10% weight), and $J_{\text{data}}$: measures structural data extraction accuracy using fuzzy matching (85% weight).
CharTeX improves upon SCRM by introducing structure-aware fuzzy matching (95% threshold) and leveraging the Hungarian algorithm for optimal alignment. In contrast to SCRM’s reliance on (entity, value) triplet matching, CharTeX incorporates column-level semantics and chrF-based scoring, enhancing robustness to text variations and structural discrepancies, particularly critical for Arabic charts with complex layouts. This design mitigates SCRM’s sensitivity to superficial mismatches and its disregard for tabular structure.
D.4 Markdown Recognition Score (MARS)
To evaluate the quality of PDF-to-Markdown conversion, we propose the Markdown Recognition Score (MARS), defined as:
$\text{MARS}=\alpha\cdot\text{chrF3}+(1-\alpha)\cdot\text{TEDS}(T_{a},T_{b})$
where $\alpha\in[0,1]$ is set to 0.5 to balance text fidelity and structural accuracy. Here, $T_{a}$ and $T_{b}$ denote the predicted and ground truth table structures, respectively.
MARS jointly captures character-level accuracy using chrF3, ideal for OCR tasks requiring fine-grained text recognition, and hierarchical layout preservation via TEDS, which quantifies the tree-edit distance between table structures. By assigning equal weight to both components, MARS offers a robust metric that reflects both semantic and structural fidelity in document conversion. As both chrF3 and TEDS are established in prior work, MARS inherits their theoretical validity without the need for further empirical justification.
D.5 Code-Oriented Diagram Metric (CODM)
The Code-Oriented Diagram Metric (CODM) extends SCRM with a graph-theoretic foundation specifically designed for diagrams where structural relationships are paramount:
$\text{CODM}=0.5J_{\text{topology}}+0.2J_{\text{topic}}+0.3J_{\text{semantics}}$ (6)
Where $J_{\text{topology}}$ evaluates diagram type (50%) using edit distance, $J_{\text{topic}}$ assesses topic identification (20%) using edit distance, and $J_{\text{semantics}}$ measures diagram structure through Graph Edit Distance (30%).
This metric converts both predicted and ground truth diagram data into graph structures, where nodes represent entities and edges represent relationships. This approach effectively evaluates both node-edge relationships and semantic labels in technical diagrams such as flowcharts, class diagrams, and sequence diagrams.
Further, domain-specific prompts are used to guide model responses for accurate metric calculation. For instance, sequence diagrams require strict adherence to Arabic UML notation standards during evaluation, ensuring fair assessment across different diagram conventions.
@@ D.5 Code-Oriented Diagram Metric (CODM)
Dataset Size GPT-4o GPT-4o-mini Azure OCR Gemini-2.0-Flash Qwen2-VL
CER WER CER WER CER WER CER WER CER WER
PATS 500 0.23 0.30 0.53 0.71 0.03 0.10 0.01 0.02 1.02 1.02
SythenAR 500 0.09 0.20 0.14 0.32 0.10 0.27 0.07 0.17 0.59 1.13
HistoryAr 200 0.51 0.82 0.67 0.96 0.24 0.68 0.28 0.64 3.46 2.86
HistoricalBooks 10 0.41 0.76 0.59 0.88 0.29 0.71 0.05 0.22 1.90 2.16
Khatt 200 0.45 0.74 0.64 0.91 0.83 0.92 0.19 0.45 1.12 5.04
Adab 200 0.30 0.73 0.35 0.83 0.99 0.99 0.19 0.56 0.63 1.08
Muharaf 200 0.56 0.90 0.63 0.94 0.52 0.82 0.33 0.69 3.57 2.87
OnlineKhatt 200 0.29 0.63 0.41 0.76 0.72 0.85 0.17 0.44 1.30 2.01
ISI-PPT 500 0.08 0.18 0.15 0.31 0.98 0.98 0.06 0.15 1.03 1.06
ArabicOCR 50 0.06 0.26 0.16 0.46 0.01 0.11 0.00 0.02 1.25 1.50
Hindawi 200 0.34 0.56 0.48 0.71 0.06 0.28 0.01 0.04 1.82 2.05
EvArest 800 0.20 0.38 0.25 0.51 0.32 0.50 0.18 0.36 0.41 0.95
3,760 0.31 0.55 0.43 0.71 0.52 0.69 0.13 0.32 1.48 1.20
Dataset Size Qwen2.5-VL AIN Qari Tesseract Surya Paddle
CER WER CER WER CER WER CER WER CER WER CER WER
PATS 500 0.98 1.03 0.26 0.36 0.00 0.00 0.14 0.28 4.66 4.67 0.77 1.00
SythenAR 500 1.68 1.69 0.21 0.40 0.04 0.16 0.31 0.72 4.82 7.90 0.80 1.01
HistoryAr 200 3.48 3.39 0.47 0.83 0.26 0.54 0.72 1.26 10.32 12.78 0.79 1.01
HistoricalBooks 10 0.67 0.97 0.33 0.72 0.84 0.88 0.74 0.99 6.81 6.30 0.71 1.00
Khatt 200 1.60 1.80 0.07 0.22 0.61 1.12 0.67 1.06 4.25 3.77 0.76 1.00
Adab 200 0.91 1.11 0.00 0.01 1.00 1.00 1.00 1.14 7.28 8.71 0.88 1.15
Muharaf 200 2.40 2.74 0.61 0.96 0.38 0.54 0.77 1.22 6.19 7.48 0.80 1.01
OnlineKhatt 200 1.52 1.53 0.36 0.70 0.03 0.12 0.59 1.20 6.71 6.95 0.78 1.03
ISI-PPT 500 1.27 1.39 0.36 0.54 0.52 0.53 0.31 0.64 4.25 3.77 0.81 1.03
ArabicOCR 50 0.02 0.08 1.00 1.00 0.01 0.01 0.01 0.01 2.75 3.58 0.77 1.00
Hindawi 200 0.27 0.42 1.00 1.00 0.11 0.15 0.31 0.72 0.15 0.20 0.76 1.00
EvArest 800 4.65 4.75 0.19 0.36 0.30 0.32 0.85 1.02 5.91 3.86 0.89 1.04
3,760 1.80 1.93 0.28 0.54 0.20 0.58 0.89 0.79 4.95 5.61 0.79 1.02